A 0.300 kg ball, initially at rest on a horizontal, frictionless surface, is struck by a 0.200 kg ball moving initially along the x axis with a speed of 2.00m/s. After the collision, the 0.200 kg ball has a speed of 1.00 m/s at an agle of = 53.0 0 to the positive x axis. (a) Determine the velocity and direction of the 0.300-kg ball after the collision
1
Expert's answer
2020-10-13T09:33:31-0400

Explanations & Calculations




  • The linear momentum of the system of 2 balls is conserved due collision in a frictionless environment.
  • After the collision the balls move away in a manner the total linear momentum is conserved.
  • Therefore, consider the momentum along the moving direction

"\qquad\qquad
\begin{aligned}
\small m_1v_1+m_2v_2&= \small m_1V_1^1+m_2V_2^{11}\\
\small 0.2\times 2+0.3\times 0&= \small 0.2\times 1\cos 53\,+ 0.3V\cos \theta\\
\small V\cos\theta&= \small 0.932ms^{-1}\cdots(1)\\
\end{aligned}"

  • That along the direction perpendicular to the moving direction

"\qquad\qquad
\begin{aligned}
\small 0.2\times 0+0.3\times 0 &= \small 0.2 \times 1\sin53\,+ 0.3\times (-V\sin\theta)\\
\small V\sin\theta &= \small 0.532 ms^{-1} \cdots(2) 
\end{aligned}"

  • Now to calculate the magnitude of the velocity "\small \to \sqrt{(1)^2+(2)^2}"

"\qquad\qquad
\begin{aligned}
\small V &= \small \sqrt{0.932^2+0.532^2}\\
&= \small \bold{1.073\,ms^{-1}}
\end{aligned}"

  • To calculate the direction"\small \to (1)/(2)"

"\qquad\qquad
\begin{aligned}
\small \tan \theta &= \small \frac{0.532}{0.932}\\
\small \theta &= \small \tan^{-1} (0.57082)\\
&= \small \bold{29.72^0}
\end{aligned}"